OpenStudy (anonymous):

another question...actually m very weak in maths...:) If Greg lost 20 pounds, then the ratio of Ted's weight to Greg's weight would be 4/3. If Ted's weighs 180 pounds what was Greg's initial weight?

OpenStudy (anonymous):

let Greg's weight be x \[\frac{180}{(x-20)}=\frac{4}{3}\] 4x-80=3*180 4x-80=540 4x=620 x=155 Greg's weight=155 pounds

OpenStudy (anonymous):

thank u sooo much....:)